OpenStudy (maddiie):

Find an equation of the line that passes through the given point and has the given slope. Express your answer in slope interpect form. (8,0),m=-3 y=-3(8)+0 ??? I'm trying here

OpenStudy (anonymous):

y=-3x+c ;now this passes thru (8,0) so 0=-24+c which gives c=24.. So the equation is y+3x=24..understood?

OpenStudy (maddiie):

Yes. i understand. What i understood was y(0)=-3(8) +c So 0= -24 because -3 x 8 = -24 +c 0=-24+c y= 3x+24